TheSmartOne (thesmartone):

Whoa, whoa! What am I missing? Calc question. Determine whether the function has a vertical asymptote or a removable discontinuity at x = −4. Graph the function using a graphing utility to confirm your answer.

TheSmartOne (thesmartone):

\(\Large f(x) = \frac{\sin (x + 4)}{x+4}\)

TheSmartOne (thesmartone):

So, if you plug in x = -4, you'll get 0 in the numerator and denominator which should make it undefined. Which should make the answer a vertical asymptote. But the correct answer is a removable discontinuity, which you can see if you graph it. https://www.desmos.com/calculator/pyx6klrxac But why is it a removable discontinuity and how is the function defined at x = -4? Also, can you show me some way to do this that wouldn't rely using a graphing calculator? I'm not sure why the homework allows you to use it. We can't use any kind of calculator on the test or quiz.

TheSmartOne (thesmartone):

\(\Large f(-4) = \sf undefined\)\[\Large \lim_{x \rightarrow -4} f(x) = 1\] This is my understanding. I don't see how f(-4) = 1 The limit as it approaches -4 can be 1, but the function cannot be defined at -4

TheSmartOne (thesmartone):

Also, on a side note, we're currently learning about continuity of a function and derivatives, so I guess the question has a relation to it (:

TheSmartOne (thesmartone):

@whpalmer4 Can you assist me? :)

zepdrix (zepdrix):

This is one of your early limit identities that you learn about,\[\large\rm \lim_{x\to0}\frac{\sin x}{x}=1\]

TheSmartOne (thesmartone):

Yeah, I learned that c:

zepdrix (zepdrix):

Think about the numerator and denominator separately,

zepdrix (zepdrix):